# Multi-Modal Data-Efficient 3D Scene Understanding for Autonomous Driving

### v1

@@ 5.4 Ablation Study
Data Splitting Strategies. For a semi-supervised learning problem, the way of partitioning labeled and unlabeled data plays a crucial. We compare four different splitting strategies (the first three are shown in Figure 4) that meet the requirements of driving data collection and perception, i.e., random, uniform, sequential sampling strategies, and the recent visual-encouraged ST-RFD [11] strategy. As shown in Tab. VIII, methods under the random or uniform sampling achieved higher results than sequential sampling. This is because the former two introduce a more diverse sample distribution than sampling sequentially. ST-RFD which picks scans using additional image cues provides a even richer labeled data set. Under all four strategies, LaserMix++ consistently achieved better performance than the baseline LaserMix [10] and MeanTeacher [18] frameworks and is competitive with Lim3D [11] on ScribbleKITTI [7].
@@ 6 Conclusion
Acknowledgments. This work is under the programme DesCartes and is supported by the National Research Foundation, Prime Minister’s Office, Singapore, under its Campus for Research Excellence and Technological Enterprise (CREATE) programme. This work is also supported by the Ministry of Education, Singapore, under its MOE AcRF Tier 2 (MOET2EP20221- 0012), NTU NAP, and under the RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s).

### v2

@@ 5.4 Ablation Study
Data Splitting Strategies. For a semi-supervised learning problem, the way of partitioning labeled and unlabeled data plays a crucial. We compare four different splitting strategies (the first three are shown in Figure 4) that meet the requirements of driving data collection and perception, i.e., random, uniform, sequential sampling strategies, and the recent visual-encouraged ST-RFD [11] strategy. As shown in Tab. VIII, methods under the random or uniform sampling achieved higher results than sequential sampling. This is because the former two introduce a more diverse sample distribution than sampling sequentially. ST-RFD which picks scans using additional image cues provides an even richer labeled data set. Under all four strategies, LaserMix++ consistently achieved better performance than the baseline LaserMix [10] and MeanTeacher [18] frameworks and is competitive with Lim3D [11] on ScribbleKITTI [7].
@@
Extension to Fully-Supervised Learning. The core principles of LaserMix, which involve partitioning and mixing LiDAR data to create diverse and challenging training samples, are not inherently limited to semi-supervised learning. These principles can be directly applied to fully-supervised settings, where the availability of labeled data allows for a more comprehensive exploration of the method’s potential. As shown in Tab. IX, our approach consistently improves by approximately 0.4% to 2.4% mIoU scores across different backbones. These results confirm that the core principles of our mixing-based methods, originally designed for semi-supervised tasks, can also significantly enhance performance in fully-supervised learning scenarios.